OpenStudy (anonymous):

Two sides of a triangle are 16 and 7. Find the size of the angle x (in radians) formed by the sides that will maximize the area of the triangle

OpenStudy (shadowlegendx):

Itʻs 1am and I am pretty much brain dead. I know the answer to this somewhere, but I canʻt think, lol

OpenStudy (shadowlegendx):

@Astrophysics I know u know tis bruthah

OpenStudy (shadowlegendx):

ty @ikram002p :)

OpenStudy (ikram002p):

like this ? 

OpenStudy (ikram002p):

i suppose yes, now area of triangle with respect to x is \(\Large ~Area ~=\frac{1}{2}~side1~\times ~side2~ \times \sin x\) as x is the angle btw side1 and side2 ok so far ?

OpenStudy (ikram002p):

so we know side1=16 and side2=7 area=1/2*7*16*sin x=56 sin x we need maximum of area, note that 0<=sin x<=1 0<=56 sin x <= 56 0<= area <= 56 which gives u maximum area of what?

OpenStudy (ikram002p):

i suppose maximum area is 56 right ? that gives us area =56 sin x 56 =56 sin x sin x =56/56=1 sin x =1 which angle x that have sin of 1 ??